OpenStudy (volleyballlover55):

Solve x^2 = 121. 60.5 −11 11 ±11

OpenStudy (volleyballlover55):

@surjithayer

OpenStudy (anonymous):

\[x^2-121=0,x^2-11^2=0 \] \[a^2-b^2=\left( a+b \right)\left( a-b \right)\]

OpenStudy (volleyballlover55):

b? -11?

OpenStudy (mahoganie.carson):

No but close

OpenStudy (mahoganie.carson):

Try again

OpenStudy (volleyballlover55):

11 c?

OpenStudy (anthonyym):

\[x^2 - 11^2 = (x+11)(x-11)\]

OpenStudy (mahoganie.carson):

hint : it has something to do with a 11

OpenStudy (anthonyym):

Or you can think of it like this: \[x^2 =121\] \[x=\pm \sqrt{11}\]

OpenStudy (volleyballlover55):

D/

OpenStudy (mahoganie.carson):

yes correct

OpenStudy (anthonyym):

Because 11*11 = 121, and -11*-11 also equals 121.

OpenStudy (volleyballlover55):

ok! :)